### v2

@@ 7 Discussion
The theoretical predictions in Fig. 7 offer indications that a central place foraging strategy is undertaken by the ants. These results suggest a temporal order in foraging, wherein ants first need to find a patch, convey the information to the rest of the colony on their way back to the nest to start patch exploitation, and then restart the search. This makes single-patch exploration-exploitation tradeoff be solved as a sequential process in time, where the exploration effort and the patch distance to the nest mainly drive the foraging efficiency. However, when there are multiple food patches, ant colonies may parallelise the exploitation as a function of the exploration effort of the colony, i.e., the number of scouts, and the initial positions from where scouting behaviour starts pol2024liquid .
@@
From a theoretical standpoint there has been a wealth of studies on the scaling of the MfFPT when the number of walkers is large and homogeneous lawley2020universal ; lawley2020probabilistic ; meerson2015mortality ; weiss1983order . Here, we have focused on experimentally relevant $\mathcal{N}$ and extended the theory to account for discrete time, the calculation of the splitting probability and also the possibility of a heterogeneous population comprised of walkers with different movement characteristics. As well as addressing the biologically motivated examples, we have utilised the theory to study some theoretical scenarios on the square lattice. We have uncovered (Fig. 4) oscillatory splitting probabilities when persistence is introduced into the population and also shown that the number of persistent walkers in the population impacts the likelihood that each target in a multi-target search is found before the other. We have confirmed, in Fig. 4(b), that similarly to the one-walker case marris2024persistent ; tejedor2012optimizing setting the persistence length too high in a multi-walker search, may cause an increase in the MfFPT (Fig. 4(b)).
@@
We note that while the mean (and global mean) first-passage times benichou2014first ; condamin2005first ; condamin2007first ; tejedor2012optimizing can be useful in quantifying the overall efficiency of searches, more recent studies have suggested that the MFPT can often over-estimate the most-probable value for the first-passage time mattos2012first and often falls in the region associated with the indirect trajectories godec2015optimization . Here, the results of Fig. 6 show clearly that the experimental times in which the food is found are associated with the direct trajectories and are more reminiscent of the modal timescale than the mean of the fFPT distribution. This corroborates with what we see experimentally where trajectories in the exploration phase tend to be direct (Fig. 1). As exemplified by these trajectories, the ants perform foraging bouts, with looping trajectories of different lengths, which do not meander and explore the entire space. Instead, these trajectories extend from the nest outwards, and then, after an unsuccessful forage, return to the nest.
@@
It is known that after the first food source is found, not only do interactions between ants increase, but the number of ants in the arena also increases cristin2024spatiotemporal . More specifically, it was observed that during the period between the first and second food source being found the number of ants in the system increased monotonically cristin2024spatiotemporal . Recognising a need to incorporate a varying number of ants into the fFPT calculations when studying foraging strategies, future work may also include the introduction of a dynamic population size into the fFPT, which is currently known only for continuous one-dimensional diffusion campos2024dynamic . Furthermore, the calculation of the fFPT to the second found food assumes the excursion from the nest to the first found patch is independent of the excursion from the first patch to the second (or back to the nest), an assumption which is clearly violated in practice. This highlights the need to develop a theory on sequential first-passage times, that is the first-passage probability to a target $\bm{s}_{2}$ conditioned on the trajectory first travelling through a previous target $\bm{s}_{1}$.
@@ Acknowledgements.
Appendix 2: The $z$-Inversion of Mean Squared Displacement
@@
The matrices are sparse as there is one unique way to move between each pair of states. For example, as seen from Eqs. (1) and (4) of the main text, the walker may never leave a positive (negative) site and find itself in a positive (negative) site in the subsequent timestep while it is in the bulk of the domain. Hence, the 12 matrices governing transitions from positive (negative) to positive (negative) sites contain only boundary interactions. Moreover, the six matrices governing the movement from states $i$ to $i$ ($\forall i\in\{1,6\}$) are identity matrices, with the sojourn probability $c$ on the main diagonal, with perturbations at the boundary elements. Finally, we have 18 matrices governing bulk movement, which each dictate one movement option.
@@
We include absorbing targets by multiplying the outgoing probability of the trap site by $(1-\rho)$, where $\rho$ is the probability of absorption. More concretely, a trap in a positive state in site $\bm{n}=(n_{1},n_{2})$ would be included by a multiplication of $(1-\rho)$ to the $(\gamma,\delta)=(n_{1},n_{2})$ element of each $\mathbb{M}(i,j)$ $(3\geq i\geq 6)$.

### v3

@@ 7 Discussion
As the theoretical predictions using the nest to patch model aligns more closely with the empirical times than the patch to patch model, (Fig. 7 and 8) there are indications that a central place foraging strategy is undertaken by the ants. These results suggest a temporal order in foraging, wherein ants first need to find a patch, convey the information to the rest of the colony on their way back to the nest, e.g. by depositing scent, to start patch exploitation, and then restart the search. In this manner the solution of the patch exploration-exploitation tradeoff is a sequential process in time, where the exploration effort and the patch distance to the nest are the main drivers to forage efficiency. However, when there are multiple food patches, ant colonies may parallelise the exploitation as a function of the exploration effort of the colony, i.e., the number of scouts, and the initial positions from where scouting behaviour starts pol2024liquid .
@@
From a theoretical standpoint there has been a wealth of studies on the scaling of the MfFPT when the number of walkers is large and homogeneous weiss1983order ; lawley2020universal ; lawley2020probabilistic ; meerson2015mortality . Here, we have focused on experimentally relevant $\mathcal{N}$ and extended the theory to account for discrete time, the calculation of the splitting probability and also the possibility of a heterogeneous population comprised of walkers with different movement characteristics. As well as addressing the biologically motivated examples, we have utilised the theory to study some theoretical scenarios on the square lattice. We have uncovered (Fig. 4) oscillatory splitting probabilities when persistence is introduced into the population and also shown that the number of persistent walkers in the population impacts the likelihood that each target in a multi-target search is found before the other. We have confirmed, in Fig. 4(b), that similarly to the one-walker case marris2024persistent ; tejedor2012optimizing setting the persistence length too high in a multi-walker search, may cause an increase in the MfFPT (Fig. 4(b)).
@@
We note that while the mean (and global mean) first-passage times tejedor2012optimizing ; benichou2014first ; condamin2005first ; condamin2007first can be useful in quantifying the overall efficiency of searches, more recent studies have suggested that the MFPT can often over-estimate the most-probable value for the first-passage time mattos2012first and often falls in the region associated with the indirect trajectories godec2015optimization . Here, the results of Fig. 6 show clearly that the experimental times in which the food is found are associated with the direct trajectories and are more reminiscent of the modal timescale than the mean of the fFPT distribution. This corroborates with what we see experimentally where trajectories in the exploration phase tend to be direct (Fig. 1). As exemplified by these trajectories, the ants perform foraging bouts, with looping trajectories of different lengths, which do not meander and explore the entire space. Instead, these trajectories extend from the nest outwards, and then, after an unsuccessful forage, return to the nest.
@@
It is known that after the first food source is found, not only do interactions between ants increase, but the number of ants in the arena also increases cristin2024spatiotemporal . More specifically, it was observed that during the period between the first and second food source being found the number of ants in the system increased monotonically cristin2024spatiotemporal . Recognising a need to incorporate a varying number of ants into the fFPT calculations when studying foraging strategies, future work may also include the introduction of a dynamic population size into the fFPT, which is currently known only for continuous one-dimensional diffusion campos2024dynamic . Furthermore, the calculation of the fFPT to the second food item assumes the excursion from the nest to the first found patch is independent of the excursion from the first patch to the second (or back to the nest), an assumption which is clearly violated in practice. This highlights the need to develop a theory on sequential first-passage times, that is the first-passage probability to a target $\bm{s}_{2}$ conditioned on the trajectory first travelling through a previous target $\bm{s}_{1}$.
@@ Acknowledgements.
Appendix 2: The $z$-Inversion of the Mean Squared Displacement
@@
The matrices are sparse as there is one unique way to move between each pair of states. For example, as seen from Eqs. (1) and (4) of the main text, the walker may never leave a positive (negative) site and find itself in a positive (negative) site in the subsequent time step when it is situated in the bulk of the domain. Hence, the 12 matrices that govern transitions from positive (negative) to positive (negative) sites contain only boundary interactions. Moreover, the six matrices governing the movement from states $i$ to $i$ ($\forall i\in\{1,6\}$) are identity matrices, with the sojourn probability $c$ on the main diagonal, with perturbations at the boundary elements. Finally, we have 18 matrices governing bulk movement, which each dictate one movement option.
@@
Finally, we include absorbing targets by multiplying the outgoing probability of the trap site by $(1-\rho)$, where $\rho$ is the probability of absorption. More concretely, a trap in a positive state in site $\bm{n}=(n_{1},n_{2})$ would be included by a multiplication of $(1-\rho)$ to the $(\gamma,\delta)=(n_{1},n_{2})$ element of each $\mathbb{M}(i,j)$ $(3\geq i\geq 6)$.